Amberjack Company is trying to decide on an allocation base to use to assign manufacturing overhead to jobs. The company has always used direct labor hours to assign manufacturing overhead to products, but it is trying to decide whether it should use a different allocation base such as direct labor dollars or machine hours.

Actual and estimated data for manufacturing overhead, direct labor cost, direct labor hours, and machine hours for the most recent fiscal year are summarized here:

Estimated Value Actual Value
Manufacturing overhead cost $ 592,000 $ 653,000
Direct labor cost $ 394,000 $ 448,000
Direct labor hours 16,300 hours 17,800 hours
Machine hours 7,300 hours 8,300 hours

Required:

1. Based on the company’s current allocation base (direct labor hours), compute the following:

a. Predetermined overhead rate. (Round your answer to 2 decimal places.)

b. Applied manufacturing overhead. (Round your intermediate calculations to 2 decimal places and final answer to the nearest whole dollar amount.)

c. Over- or underapplied manufacturing overhead. (Round your intermediate calculations to 2 decimal places and final answer to the nearest whole dollar amount.)

2. If the company had used direct labor dollars (instead of direct labor hours) as its allocation base, compute the following:

a. Predetermined overhead rate. (Round your answer to 2 decimal places, i.e. 3.63%)

b. Applied manufacturing overhead. (Round your intermediate calculations to 2 decimal places and final answer to the nearest whole dollar amount.)

c. Over- or underapplied manufacturing overhead. (Round your intermediate calculations to 2 decimal places and final answer to the nearest whole dollar amount.)

3. If the company had used machine hours (instead of direct labor hours) as its allocation base, compute the following:

a. Predetermined overhead rate. (Round your answer to 2 decimal places.)

b. Applied manufacturing overhead. (Round your intermediate calculations to 2 decimal places and final answer to the nearest whole dollar amount.)

c. Over- or underapplied manufacturing overhead. (Round your intermediate calculations to 2 decimal places and final answer to the nearest whole dollar amount.)

4. Based on last year’s data alone, which allocation base would have provided the most accurate measure for applying manufacturing overhead costs to production?
Answer #1

ANSWER

1a) Overhead rate = Estimated overhead/Estimated labor hours = 592,000/16,300 = $36.32 per labor hour

1b) Applied overhead = 17,800*36.32 = $646,496

1c) Over or under applied manufacturing overhead = 653,000-646,496= 6,504 Under applied

2a) Overhead rate = 592,000/394,000 = 150.25%

2b) Applied overhead = 448,000*150.25% = $673,120

2c) Over or under applied overhead = 653,000-673,120= $20,120 Over applied

3a) Overhead rate = 592,000/7,300 = $81.09 per machine hour

3b) Applied overhead = 8,300*81.09 = $673,047

3c) Over or under applied overhead = 673,047-653,000 = $20,047 Over applied

__________________________________

If you have any query or any Explanation please ask me in the comment box, i am here to helps you.please give me positive rating.

*****************THANK YOU*************